OpenStudy (anonymous):

what's the expanded form of x²+5x?

OpenStudy (anonymous):

you first take out what's common: x(x+5) = 0 ^_^

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

see i tried this on the graphing calculator thing and it doesn't show up with the graph i had for the original equation

OpenStudy (anonymous):

how come?

OpenStudy (anonymous):

Did you try and draw a graph x(x+5) = 0? Because that would only have lines at x = 0, -5. You'd need y or f(x) or whatever you use for that axis = x(x+5)

OpenStudy (anonymous):

well like i am working on this problem- a square has side of lengths x centimeters. A new rectangle is created by increasing one dimension by 5cm. So the factored form is x^2+5x, but I need toe expanded to match it too

OpenStudy (anonymous):

Actually, x(x+5) is the factored form, and x^2 + 5x is the expanded form.

OpenStudy (anonymous):

oh okay thanks